OpenStudy (wintersuntime):

why are there 4 answers to the imaginary number i? The 4 answers are -1, -i, 1, and j. Can someone help me ? this is 10th grade math

OpenStudy (irishboy123):

maybe you are looking at the 4 fourth roots.

OpenStudy (wintersuntime):

okay thank you

OpenStudy (anonymous):

i think you are looking at \(i\) to various powers

OpenStudy (anonymous):

not the fourth roots, which is completely different

OpenStudy (anonymous):

\[i^1=i\\ i^2=-1\\ i^3=-i\\ i^4=1\]

OpenStudy (wintersuntime):

oh okay thank you that makes more sense

OpenStudy (fibonaccichick666):

I will also add, that i is just a defined variable. i always is equal to the square root of negative 1. That is its definition. Then I agree with sat, I think that is what you were actually looking at

OpenStudy (wintersuntime):

thank you but why are the various powers to i equal 4 same answers

OpenStudy (anonymous):

\[i^2=-1\] is the definition of \(i\) the number whose square is \(-1\) i.e \(i=\sqrt{-1}\)

OpenStudy (anonymous):

so \(i^3=-1\times i=-i\)

OpenStudy (fibonaccichick666):

ah, so i like I said, is the square root of -1, now. If you take \[\sqrt{-1}*\sqrt{-1}=(\sqrt{-1})^2=-1\]

OpenStudy (anonymous):

and \(i^4=-i\times i=-(-1)=1\)

OpenStudy (fibonaccichick666):

then sats, for cube and fourth

OpenStudy (anonymous):

once you get to \(i^4=1\) it should be pretty clear that you start over

OpenStudy (wintersuntime):

oh alright thank you guys

OpenStudy (anonymous):

\[i^4=1\\ i^5=1\times i=i\\ i^6=i\times i=-1\\ etc\]

OpenStudy (anonymous):

\[i^{2003}=i^3=-i\]

OpenStudy (fibonaccichick666):

np